OpenStudy (anonymous):

Find the set of values taken by 1\(9x^2+12x+7) for real values of X . * In the first part of the question they asked me to make out the complete square of the quadratic formula alone with being under 1 . It was (3x+2)^2+3 .

OpenStudy (ikram002p):

9x^2+12x+7 \(\neq\) 0 find value that makes it 0 , then domain would be R-|zeros of it |

OpenStudy (anonymous):

hey is the answer is to be given in the perspective of calculus

OpenStudy (anonymous):

I guess that basically the book is asking the range of \[f(x)= \frac{ 1 }{ 9x^2 + 12x + 7 }\] WHat's yar book answer?

OpenStudy (anonymous):

@Lama97

OpenStudy (ikram002p):

oh i thought it wont the domain :P

OpenStudy (anonymous):

its 0<f(x) , or equal 1\3

OpenStudy (anonymous):

ikram I just thought the same way

OpenStudy (anonymous):

oh i meant less or equal to 1\3

OpenStudy (ikram002p):

:) then u need to find max which would be 1/3 and note that the function >0 , since 9x^2+12x+7 is positive

OpenStudy (ikram002p):

do you know how to find max ?

OpenStudy (anonymous):

max point ? vertex right ?

OpenStudy (ikram002p):

first find derivative of the function , f'(x) maximum point or minimum point would be when f'(x) =0

OpenStudy (anonymous):

I think that it wants the possible solution for all the x on the no. line SO therefore we have to prove that f(x) > 0 So we have to fing the maximma(or maximum value of f(x0 So I guess to solve this we have to differentiate the EQN. SO let : \[h(x) = 9x^2+12x+7\] y = h(x) SO : We know that: \[\frac{ d }{ dx }x^n = nx^{n-1}\] \[\frac{ d }{ dx }h(x) = 18x+12 \] Then h'(x) = 18x + 12(which is at min or max when h'(x) = 0) SO: If 18x+12 = 0 then x =-2/3(This is minimma I guess) Therefore I think that this point becomes a maximum in f(x)('cuz f(x) = 1/h(x)). Therefore the maximum of f(x) is f(-2/3) = 1/3 (just plug the number in the f(x)..Lol) Isn't g(x) is at a minimum for x=-2/3 and has the value of 3 and all other values of g(x) are greater than this. SO this means that f(x) is always +ve but >1/3 (>0 and=0 at infinity) SO the answer is: \[f(x)> 0\] Or \[f(x) = 1/3\]

OpenStudy (ikram002p):

f(-2/3)=1/3 :D

OpenStudy (anonymous):

@Lama97

OpenStudy (anonymous):

guys you both did a great job , to whom should I give the medal :P

OpenStudy (ikram002p):

to @Adjax :) he explained it :P i only gives hint xD

OpenStudy (anonymous):

well yes he explained it in details. but you both really deserve it. thanks ikram :) .. Adjax my great appreciation :)